1. A 0.75-m3 rigid tank initially contains air whose density is 1.18 kg/m3. The tank is connected to a high-pressure supply line through a valve. The valve is opened, and air is allowed to enter the tank until the density in the tank rises to 4.95 kg/m3. Determine the mass of air that has entered the tank. Answer: 2.83 kg


2. Consider the flow of an incompressible Newtonian fluid between two parallel plates that are 4 mm apart. If the upper plate moves to the right with u1 = 5 m/s while the bottom one moves to the left with u2 = 1.5 m/s, what would be the net flow rate at a cross-section between two plates? Take the plate width to be 5 cm.


1
Expert's answer
2021-12-14T09:28:36-0500

1.

"\Delta m=\Delta\rho V=2.83~kg,"

2.

"Q=VS=\frac{v_2-v_1}2mn=3.85\cdot 10^{-5}~\frac{m^3}s."